OpenStudy (anonymous):

Calculate the wavelength, in nanometers, of the spectral line produced when an electron in a hydrogen atom undergoes the transition from the energy level n = 7 to the level n = 2.

OpenStudy (anonymous):

Rydberg formula \[ \frac{1}{\lambda} = R \Bigg( \frac{1}{n_1^2}-\frac{1}{n_2^2}\Bigg)\] \[n_1=2 \quad ; \quad n_2=7\] \[R=1.0973 \times 10^7 \ m^{-1}\]